Chlorine can be prepared in the laboratory by the reaction of manganese dioxide with hydrochloric acid, HCl(aq)HCl(aq), as described by the chemical equation

MnO2(s)+4HCl(aq)⟶MnCl2(aq)+2H2O(l)+Cl2(g)MnO2(s)+4HCl(aq)⟶MnCl2(aq)+2H2O(l)+Cl2(g)

How much MnO2(s)MnO2(s) should be added to excess HCl(aq)HCl(aq) to obtain 385 mL Cl2(g)385 mL Cl2(g) at 25 °C and 805 Torr805 Torr?

When aluminum is placed in concentrated hydrochloric acid, hydrogen gas is produced.

2Al(s)+6HCl(aq)⟶2AlCl3(aq)+3H2(g)2Al(s)+6HCl(aq)⟶2AlCl3(aq)+3H2(g)

What volume of H2(g)H2(g) is produced when 2.202.20 g Al(s)Al(s) reacts at STP?

Hydrazine, N2H4N2H4, reacts with oxygen to form nitrogen gas and water.

N2H4(aq)+O2(g)⟶N2(g)+2H2O(l)N2H4(aq)+O2(g)⟶N2(g)+2H2O(l)

If 2.45 g2.45 g of N2H4N2H4 reacts with excess oxygen and produces 0.850 L0.850 L of N2N2, at 295 K295 K and 1.00 atm,1.00 atm, what is the percent yield of the reaction?

The average kinetic energy of the molecules in a gas sample depends only on the temperature, ?.T. However, given the same kinetic energies, a lighter molecule will move faster than a heavier molecule, as shown in the equation for rms speed

rms speed=3??ℳ⎯⎯⎯⎯⎯⎯⎯⎯⎯√rms speed=3RTℳ

where ?=8.314 J/(mol⋅K)R=8.314 J/(mol⋅K) and ℳℳ is molar mass in kilograms per mole. Note that a joule is the same as a kilogram‑meter squared per second squared (kg·m2/s2).

What is the rms speed of N2N2 molecules at 299 K?299 K?

rms speed:

m/sm/s

What is the rms speed of HeHe atoms at 299299 K?

rms speed:
Answer #1

1.

Using ideal gas equation

PV = nRT

Or, number of moles of Cl2 gas ( n ) =

Given , pressure(P) = 895 Torr = 895 torr ×  = 1.17 atm

Volume (V) = 385 mL = 385mL × = 0.385 L.

Temperature (T) = 250c = 298 K.

R = 0.082 L-atm/mol.K

Now, n = = 0.018 .

From the balanced reaction

Mole ratio of MnO2 and Cl2 = 1:1 .

Hence moles of MnO2 = 0.018

Mass of MnO2 should be added

= moles* of MnO2 * molar mass of MnO2

= 0.018 * 87 = 1.556 g.

2.

Moles of Al = (mass of Al/molar mass of Al)

= ( 2.20/27) = 0.0814

From balanced equation

Mole ratio of Al and H2 = 2:3

Then, moles of H2 gas = 0.0814 × = 0 .122

Mass of hydrogen gas = ( 0.122*2) = 0.244 g.

[ Molar mass of hydrogen = 2 g/mol]

3.

Moles of N2H4 = (2.45/32) = 0.0765

From balanced equation :

Moles of N2 produced = moles of N2H4 = 0.0765

Then volume of N2 = (nRT/P) = (0.0765*0.082*295/1)

= 1.850 L

Percent yield

= =  

= 45.94%.